Does anyone know the answer to this problem?
The sum of three consecutive odd integers is 201. I need to find the integers.
Thank you
Adriana

Answer by tutorcecilia(2144)

The sum: addition or "added together"
three consecutive odd integers: x+2 and x+4
is 201: =201
(x+2)+(x+4)=201 [Simplify]
x+2+x+4=201
2x+6=201 [Solve for x]
2x+6-6=201-6
2x=195
2x/2=195/2
x=97.5 +2 = 99.5 is the first integer
x+1 or 97.5+4=101.5 is the second integer
Therefore 99.5+101.5=201
.
Check by plugging (x=97.5) back into the original equation:
(x+2)+(x+4)=201
(97.5+2)+(97.5+4)=201
201=201 [Checks out]
.
If the problem ask for "consecutive numbers" use:
x=the first consecutive number
x+1=the second consecutive number
x+3=the third consecutive number
.
If the problem ask for consecutive "odd" or consecutive "even" numbers use:
x+2=the first consecutive number
x+4=the second consecutive number
x+6= the third consecutive number, etc.